# Multi-bandpass Photometry for Exoplanet Atmosphere Reconnaissance (MPEAR) with the Habitable Worlds Observatory (HWO) I. Differentiating Earth from Neptunes During Discovery

### v1

@@ 7.2 $S/N$ Considerations
Investing more time in single-epoch observations compared to multiple visits might prove to be useful in the observing strategy for HWO. Achieving definitive differentiation would require S/NS/N=15-20, which would translate in 6-10 hours of observing time. While this doubles or triples the observing time compared to the baseline 3.2 hours required for S/NS/N=7 (see Section 5), it is still substantially less than spectroscopic alternatives and avoids the orbital confusion that appears in multi-epoch strategies. It would also be a time-saver in terms of overheads: slew, dark hole digging, thermal stabilization upon pointing would only be performed once for a single epoch, while they would need to be repeated for multiple visits. However, these findings will need be further assessed in the context of the overall yield of the mission (see Section 7.4).
7.3 Comparing with Previous Work
@@
JKT2016 answered a similar question to what we aim to do in this study. Through maximization of the distance between the Earth and any other planet in color-color space, the authors could identify combinations of bandpasses that would allow differentiation of the Earth from false positives and found optimized bandpasses around 431-531 nm, 569-693 nm, 770-894 nm. These values stem from the optimization algorithm that they used, spanning all possible wavelength combinations and do not take into account any noise associated with the measured flux to identify the optimized bandpasses. Rule-of-thumb calculations on the desired noise level to differentiate the Earth from other false positives was considered at a later stage, where the authors discussed the value of photometry compared to spectroscopy in detector-dominated regimes. While this is outside the scope of this work (see Section 7.4), a critical advancement in our approach is the integration of realistic noise modeling throughout the bandpass optimization process.
@@
The color-color space distance is a metric that we also consider in this paper (see Section 6), although we identify the best bandpasses based on their differentiation potential, considering the expected noise for each photometric band (see Section 5). We therefore include the noise at an earlier stage in the process and that informs the bandpass optimization directly. This can be seen for example in the near-infrared, where the significance decreases as the noise increases and no bandpasses are chosen at longer wavelengths, although noticeable differences between the Earth and the cold Neptune could be identified (see Figure 5).
@@
JKT2016 considered a wider range of planets for their calculation, though phase was mostly kept fixed at 60 except for some cases. In our work, we consider phases that could be reasonably plausible with a given detection (satisfying the orbital condition, see Section 3). Results are however highly dependent on phase, and we should expect variations when assuming an Earth-like planet not at quadrature, as was highlighted in JKT2016.
@@
7.4 Limitations and Future Work
@@
We focused on differentiating the Earth from (warm and cold) Neptune-like planets since it was motivated by previous work showing how these planet archetypes could show similar features (2018haex.bookE..98R). Even in this simplified case, we could already demonstrate how bandpass optimization is extremely dependent on specific planetary spectral features. We can imagine that, extending the study over multiple planetary classes covering a variety of potential atmospheres, would likely yield different bandpass optimizations. We also did not consider planet occurrence rates in our analysis, though they might play a role in determining how likely it might be that an Earth-like planet might be confused with a Neptune on a specific orbit. Occurrence rates for HZ terrestrial planets around Solar-like stars (as well as larger planets in more distant orbits) are still lacking, and are expected to be one of the main outcomes of upcoming missions such as PLATO (PLAnetary Transits and Oscillations of stars Rauer2024-PLATOMission). Studies like this could be periodically carried out as these estimates are refined.
@@
We employed fixed templates for the warm and cold Neptune, allowing the flux to only be impacted by orbital parameters and focusing our attention on the effects of orbital geometry on discrimination performance. We expect, however, that the atmosphere of a planet would also be impacted by its orbital features: the instellation received varies dramatically with distance from the star and that would impact both the thermal profile and the atmospheric chemistry. A more accurate approach would be to calculate a grid of models that takes into account variations in atmospheric profiles, although an accurate computation would require a larger effort and substantial computing power. The nature of such atmospheres is also predicted to be very different: gaseous planets span a range of cloud pressures and methane and water vapor mixing ratios above the cloud deck, all of which impacts the spectral features and its shape. The radius could also be varied, for example smaller radii for warm Neptunes which are predicted to be abundant and likely to cause confusion. Furthermore, we have not considered variations in the spectrum of the Earth during its evolution, nor the broader range of terrestrial atmospheres that have been explored in earlier studies (JKT2016; Smith2020). Future studies are foreseen to explore different planetary classes to provide statistically robust bandpass combinations that could separate habitable planets from false positives. This could be done with the methodology developed in this study, which is fully independent from the underlying physics that shape the spectra.
@@
In this work, we assumed circular orbits, prescribing a planet detected at a specific projected separation. In reality, orbits can be eccentric and their orientation can also vary ($\Omega$ is not fixed at $0^{\circ}$). The effect of this assumption is negligible in this work, since we are dealing with a single epoch and the apparent phase only depends from the angle between the star and the planet, and the one between the planet and the observer (see Section 3). However, this would play a decisive role when considering multiple epochs or when considering the climate evolution of planets in eccentric orbits. We also assumed that no other known planets are in the system; in reality, a known wider companion might provide information on the likely orientation of the orbital plane, which could reduce orbital degeneracies and improve discrimination strategies. In a future publication we expect to address the added benefit of multi-bandpass photometry as a qualitative discrimination method to correctly associate observations of multi-planetary systems performed at different epochs, which could simplify the orbit disentanglement.
@@ 8 Summary
We developed a comprehensive algorithm that explores the parameter space and identifies potentially confusing scenarios, both in terms of orbital parameters and flux in the discovery bandpass. We modeled orbits of potential (cold and warm) Neptune-like planets that would be consistent with the same on-sky projected distance as the Earth at quadrature orbiting a Sun-twin star at 10 pc. We then modeled the corresponding spectra and found the scenarios that would yield a comparable flux at 500 nm (the discovery bandpass assumed by HWO), thus making the Earth and the Neptune-like planets virtually impossible to distinguish with a single photometric data point. We then evaluated the noise that any other potential bandpass across the wavelength range would have for the exposure time determined by the discovery observation (to achieve S/NS/N=7, approximately 3.2 hours). To achieve this, we used the HWO coronagraph exposure time calculator that we developed, which we call pyEDITH. We then identified the 2- and 3-bandpass combinations that would yield the largest statistical significance in differentiating between the spectrum of an Earth and those of warm and cold Neptune-like planets, i. e., the largest differentiation potential to distinguish an Earth from a Neptune. We find that:
@@
The authors thank Rhonda Morgan, Samantha Hasler, Sarah Steiger, Natasha Latouf, and Corey Spohn for helpful discussions. The work of E.A. and M.H.C. was supported by appointments to the NASA Postdoctoral Program at the NASA Goddard Space Flight Center, administered by Oak Ridge Associated Universities under contract with NASA (ORAU-80HQTR21CA005). A.M.M. and A.V.Y acknowledge support from the GSFC Sellers Exoplanet Environments Collaboration (SEEC), which is supported by NASA’s Planetary Science Division’s Research Program. A.M.M. acknowledges support from the Explanet Spectroscopy Technology ISFM, which is supported by NASA’s Astrophysics Science Division. {contribution}E.A. developed the MPEAR algorithm, developed and performed the necessary pyEDITH software enhancements, performed the analysis, wrote and edited the manuscript. E.A., A.M.M., and A.R. conceptualized the project. A.M.M., A.R. and C.C.S. provided advice. M.H.C. co-developed pyEDITH. A.P., V.K., G.L.V., and R.H. provided input templates. V.K., A.P., and A.V.Y. wrote portions of the manuscript. All listed authors reviewed and edited the manuscript.

### v2

@@ VII.2 $S/N$ Considerations
Investing more time in single-epoch observations compared to multiple visits might prove to be useful in the observing strategy for HWO. Achieving definitive differentiation would require S/NS/N=15-20, which would translate in 6-10 hours of observing time. While this doubles or triples the observing time compared to the baseline 3.2 hours required for S/NS/N=7 (see Section V), it is still substantially less than spectroscopic alternatives and avoids the orbital confusion that appears in multi-epoch strategies. It would also be a time-saver in terms of overheads: slew, dark hole digging, thermal stabilization upon pointing would only be performed once for a single epoch, while they would need to be repeated for multiple visits. However, these findings will need be further assessed in the context of the overall yield of the mission (see Section VII.4).
VII.3 Comparing with Previous Work
@@
Krissansen-Totton et al. (2016) answered a similar question to what we aim to do in this study. Through maximization of the distance between the Earth and any other planet in color-color space, the authors could identify combinations of bandpasses that would allow differentiation of the Earth from false positives and found optimized bandpasses around 431-531 nm, 569-693 nm, 770-894 nm. These values stem from the optimization algorithm that they used, spanning all possible wavelength combinations and do not take into account any noise associated with the measured flux to identify the optimized bandpasses. Rule-of-thumb calculations on the desired noise level to differentiate the Earth from other false positives was considered at a later stage, where the authors discussed the value of photometry compared to spectroscopy in detector-dominated regimes. While this is outside the scope of this work (see Section VII.4), a critical advancement in our approach is the integration of realistic noise modeling throughout the bandpass optimization process.
@@
The color-color space distance is a metric that we also consider in this paper (see Section VI), although we identify the best bandpasses based on their differentiation potential, considering the expected noise for each photometric band (see Section V). We therefore include the noise at an earlier stage in the process and that informs the bandpass optimization directly. This can be seen for example in the near-infrared, where the significance decreases as the noise increases and no bandpasses are chosen at longer wavelengths, although noticeable differences between the Earth and the cold Neptune could be identified (see Figure 5).
@@
Krissansen-Totton et al. (2016) considered a wider range of planets for their calculation, though phase was mostly kept fixed at 60 except for some cases. In our work, we consider phases that could be reasonably plausible with a given detection (satisfying the orbital condition, see Section III). Results are however highly dependent on phase, and we should expect variations when assuming an Earth-like planet not at quadrature, as was highlighted in Krissansen-Totton et al. (2016).
@@
VII.4 Limitations and Future Work
@@
We focused on differentiating the Earth from (warm and cold) Neptune-like planets since it was motivated by previous work showing how these planet archetypes could show similar features (Roberge and Seager, 2018). Even in this simplified case, we could already demonstrate how bandpass optimization is extremely dependent on specific planetary spectral features. We can imagine that, extending the study over multiple planetary classes covering a variety of potential atmospheres, would likely yield different bandpass optimizations. We also did not consider planet occurrence rates in our analysis, though they might play a role in determining how likely it might be that an Earth-like planet might be confused with a Neptune on a specific orbit. Occurrence rates for HZ terrestrial planets around Solar-like stars (as well as larger planets in more distant orbits) are still lacking, and are expected to be one of the main outcomes of upcoming missions such as PLATO (PLAnetary Transits and Oscillations of stars Rauer et al., 2024). Studies like this could be periodically carried out as these estimates are refined.
@@
We employed fixed templates for the warm and cold Neptune, allowing the flux to only be impacted by orbital parameters and focusing our attention on the effects of orbital geometry on discrimination performance. We expect, however, that the atmosphere of a planet would also be impacted by its orbital features: the instellation received varies dramatically with distance from the star and that would impact both the thermal profile and the atmospheric chemistry. A more accurate approach would be to calculate a grid of models that takes into account variations in atmospheric profiles, although an accurate computation would require a larger effort and substantial computing power. The nature of such atmospheres is also predicted to be very different: gaseous planets span a range of cloud pressures and methane and water vapor mixing ratios above the cloud deck, all of which impacts the spectral features and its shape. The radius could also be varied, for example smaller radii for warm Neptunes which are predicted to be abundant and likely to cause confusion. Furthermore, we have not considered variations in the spectrum of the Earth during its evolution, nor the broader range of terrestrial atmospheres that have been explored in earlier studies (Krissansen-Totton et al., 2016; Smith et al., 2020). Future studies are foreseen to explore different planetary classes to provide statistically robust bandpass combinations that could separate habitable planets from false positives. This could be done with the methodology developed in this study, which is fully independent from the underlying physics that shape the spectra.
@@
In this work, we assumed circular orbits, prescribing a planet detected at a specific projected separation. In reality, orbits can be eccentric and their orientation can also vary ($\Omega$ is not fixed at $0^{\circ}$). The effect of this assumption is negligible in this work, since we are dealing with a single epoch and the apparent phase only depends from the angle between the star and the planet, and the one between the planet and the observer (see Section III). However, this would play a decisive role when considering multiple epochs or when considering the climate evolution of planets in eccentric orbits. We also assumed that no other known planets are in the system; in reality, a known wider companion might provide information on the likely orientation of the orbital plane, which could reduce orbital degeneracies and improve discrimination strategies. In a future publication we expect to address the added benefit of multi-bandpass photometry as a qualitative discrimination method to correctly associate observations of multi-planetary systems performed at different epochs, which could simplify the orbit disentanglement.
@@ VIII Summary
We developed a comprehensive open-source algorithm [^9] Available at: https://github.com/eleonoraalei/mpear that explores the parameter space and identifies potentially confusing scenarios, both in terms of orbital parameters and flux in the discovery bandpass (Alei, 2026) . We modeled orbits of potential (cold and warm) Neptune-like planets that would be consistent with the same on-sky projected distance as the Earth at quadrature orbiting a Sun-twin star at 10 pc. We then modeled the corresponding spectra and found the scenarios that would yield a comparable flux at 500 nm (the discovery bandpass assumed by HWO), thus making the Earth and the Neptune-like planets virtually impossible to distinguish with a single photometric data point. We then evaluated the noise that any other potential bandpass across the wavelength range would have for the exposure time determined by the discovery observation (to achieve S/NS/N=7, approximately 3.2 hours). To achieve this, we used the HWO coronagraph exposure time calculator that we developed, which we call pyEDITH. We then identified the 2- and 3-bandpass combinations that would yield the largest statistical significance in differentiating between the spectrum of an Earth and those of warm and cold Neptune-like planets, i. e., the largest differentiation potential to distinguish an Earth from a Neptune. We find that:
@@
The authors thank Rhonda Morgan, Samantha Hasler, Sarah Steiger, Natasha Latouf, and Corey Spohn for helpful discussions. The work of E.A. and M.H.C. was supported by appointments to the NASA Postdoctoral Program at the NASA Goddard Space Flight Center, administered by Oak Ridge Associated Universities under contract with NASA (ORAU-80HQTR21CA005). A.M.M. and A.V.Y acknowledge support from the GSFC Sellers Exoplanet Environments Collaboration (SEEC), which is supported by NASA’s Planetary Science Division’s Research Program. A.M.M. acknowledges support from the Explanet Spectroscopy Technology ISFM, which is supported by NASA’s Astrophysics Science Division.
{contribution}
E.A. developed the MPEAR algorithm (released as open-source software (Alei, 2026)), developed and performed the necessary pyEDITH software enhancements, performed the analysis, wrote and edited the manuscript. E.A., A.M.M., and A.R. conceptualized the project. A.M.M., A.R. and C.C.S. provided advice. M.H.C. co-developed pyEDITH. A.P., V.K., G.L.V., and R.H. provided input templates. V.K., A.P., and A.V.Y. wrote portions of the manuscript. All listed authors reviewed and edited the manuscript.